3) EagleShare is a popular new ridesharing app. Usage has been increasing rapidly, and the founders have realized that they need to upgrade their data center, as their servers will soon be unable to handle the level of activity. The metric EagleShare uses to measure activity is the number of “requests.” A request occurs each time a passenger or driver searches for a ride, accepts or rejects a ride, and completes a ride. Currently, they are averaging 10,000 requests per day. The average marginal profit associated with a request (ignoring any server costs) is $0.20.

Since they are not IT architecture experts, they have decided to outsource hosting and management of their servers. They are considering three possible 1-year contracts. The server usage involved with each contract has been converted into number of requests.

Contract A costs $20,000, and includes 400,000 requests.
Contract B costs $50,000, and includes 1,200,000 requests.
Contract C costs $100,000, and includes 3,000,000 requests.

For the sake of simplicity, assume that when a contract’s capacity is reached, app users will temporarily switch to a competitor; EagleShare will receive no more profit for the year, but will not suffer any additional goodwill cost or loss of future revenue.

NOTE: SHOW ALL YOUR WORK. USE 4 DECIMAL PLACES IN YOUR CALCULATIONS.

  1. (1 point) If EagleShare were not anticipating any growth, and expected the average rate of 10,000 requests per day to continue for the next year, which contract should they choose?

  1. (3 points) Compute the breakeven point (in # of requests) for each of the three contracts.

  1. (1 point) For what # of requests would the profit with Contract A and with Contract B be equal?
  1. (1 point) For what # of requests would the profit with Contract B and with Contract C be equal?
  1. (1 point) If expected demand for the next year is 1,000,000 requests, what is the capacity cushion with Contract B?
Answer #1

a. If EagleShare expects the average rate of 10000 requests per day to continue for the next year, then it should choose the contract with the lowest cost which can cater a request volume of minimum 10000. We can see that contract A is the contract with lowest cost which can cater to 10000 request. Hence Contract A should be chosen.

b. Breakeven point is given by Contract Cost divided by marginal profit.

For contract A, Breakeven point = 20,000/0.2 = 100,000 requests

For contract B, Breakeven point = 50,000/0.2 = 250,000 requests

For contract C, Breakeven point = 100,000/0.2 = 500,000 requests

c. As we can see, the cost of Contract B is more than cost of Contract A by 30,000. As there is no marginal cost involved till the time requests are 400,000 contract A will always have 30,000 more in profit than contract B. Once the requests exceed 400,000 contract A would not cater to any additional requests and hence profit from Contract A will stay the same. However Contract B can still cater to additional request. The profit from both the contracts will be equal when the number of additional requests above 400,000 will give a profit which is equal to 30,000 which is difference in costs of 2 contracts. As each request gives a profit of 0.2, the number of additional requests required will be 30,000/0.2 = 150,000. Hence the total number of requests is 400,000 + 150,000 = 550,000. Thus when requests are 550,000 profit from contract A and B would be equal.

d. As we can see, the cost of Contract C is more than cost of Contract B by 50,000. As there is no marginal cost involved till the time requests are 1,200,000 contract B will always have 50,000 more in profit than contract C. Once the requests exceed 1,200,000 contract B would not cater to any additional requests and hence profit from Contract B will stay the same. However Contract C can still cater to additional request. The profit from both the contracts will be equal when the number of additional requests above 1,200,000 will give a profit which is equal to 50,000 which is difference in costs of 2 contracts. As each request gives a profit of 0.2, the number of additional requests required will be 50,000/0.2 = 250,000. Hence the total number of requests is 1,200,000 + 250,000= 1 450,000. Thus when requests are 1,450,000 profit from contract B and C would be equal.

e. Capacity Cushion is the difference between utilized capacity and available capacity. As available capacity for Contract B is 1,200,000 and utilized capacity is expected to be 1,000,000, capacity cushion = 1,200,000-1,000,000 = 200,000 requests.